# Generative Modeling with Bayesian Sample Inference

### v2

@@ Abstract
We derive a novel generative model from iterative Gaussian posterior inference. By treating the generated sample as an unknown variable, we can formulate the sampling process in the language of Bayesian probability. Our model uses a sequence of prediction and posterior update steps to iteratively narrow down the unknown sample starting from a broad initial belief. In addition to a rigorous theoretical analysis, we establish a connection between our model and diffusion models and show that it includes Bayesian Flow Networks (BFNs) as a special case. In our experiments, we demonstrate that our model improves sample quality on ImageNet32 over both BFNs and the closely related Variational Diffusion Models, while achieving equivalent log-likelihoods on ImageNet32 and CIFAR10.
@@ 1 Introduction
The field of deep learning has produced a multitude of generative models over the years (Harshvardhan et al., 2020). Variational autoencoders, for example, learn the data distribution by compressing data into a lower-dimensional representation (Kingma & Welling, 2013). Normalizing flows learn to map between a prior and the data distribution via invertible transformations, enabling exact likelihood computation (Rezende & Mohamed, 2015). Generative adversarial networks generate samples by pitting two models against each other such that one proposes artificial data samples while the other tries to distinguish real and generated (Goodfellow et al., 2014). Recently, diffusion models (DMs) have become a cornerstone of generative modeling (Sohl-Dickstein et al., 2015; Ho et al., 2020). They define a multi-step forward process that gradually adds noise to the data, turning it into pure noise. Then, a model is trained to reverse this process, enabling the generation of new data samples by starting from noise and iteratively denoising.
@@ 3 Sample Generation with Posterior Inference
0: Initial precision $\lambda_{0}$,
precision schedule $\alpha_{i}$ for $i\in[k]$
@@
Algorithm 1 Sampling with posterior inference
Let $f_{{\bm{\theta}}}:\mathbb{R}^{n}\times\mathbb{R}_{+}\to\mathbb{R}^{n}$ be a learned model with parameters $\bm{\theta}$} that estimates which unknown sample $\bm{x}$} we have observed so far from our current belief $({\bm{\mu}}_{i},\lambda_{i})$. We estimate $\bm{x}$} as $\hat{{\bm{x}}}_{i-1}=f_{{\bm{\theta}}}({\bm{\mu}}_{i-1},\lambda_{i-1})$ and sample a noisy measurement ${\bm{y}}_{i}\sim\mathcal{N}_{\mathrm{P}}(\hat{{\bm{x}}}_{i-1},\alpha_{i})$ of $\hat{{\bm{x}}}_{i-1}$ in place of $\bm{x}$} with precision $\alpha_{i}$. Then, we can update our belief with ${\bm{y}}_{i}$ and Lemma˜2.1 to the posterior $({\bm{\mu}}_{i},\lambda_{i})$. Now, we alternate between these two steps, i.e. predicting and taking a noisy measurement followed by updating our current belief, until the posterior precision $\lambda_{i}$ is sufficient. Finally, we return $\hat{{\bm{x}}}^{*}=f_{{\bm{\theta}}}({\bm{\mu}}_{k},\lambda_{k})$ as our generated sample. See ˜9 for a formal description and Fig.˜1 for a visual explanation.
@@ 3.1 Evidence Lower Bound
By interpreting BSI as a hierarchical latent variable model, we derive an ELBO (Kingma & Welling, 2013), i.e. a lower bound on $\log\mathrm{p}({\bm{x}})$ assigned to a data point by our model. The ELBO will then serve as a natural training target for $f_{{\bm{\theta}}}$ to ensure that true data samples have high likelihood under our model.
@@
$\mathrm{p}({\bm{x}})=\operatorname*{\mathbb{E}}_{\mathrm{p}({\bm{\mu}}_{0})\cdot\mathrm{p}({\bm{\mu}}_{1}\mid\mathrm{p}({\bm{\mu}}_{0})\cdots\mathrm{p}({\bm{\mu}}_{k}\mid{\bm{\mu}}_{k-1}))}[\mathrm{p}({\bm{x}}\mid{\bm{\mu}}_{k})].$ (2)
@@
$\mathcal{L}_{\mathrm{R}}=\operatorname*{\mathbb{E}}_{\mathrm{q}({\bm{\mu}}_{k}\mid{\bm{x}},\lambda_{k})}\big[{\shortminus}\negthinspace\log\mathcal{N}_{\mathrm{P}}({\bm{x}}\mid\hat{{\bm{x}}}_{k},\alpha_{\mathrm{R}})\big]\quad\text{and}\quad\mathcal{L}_{\mathrm{M}}^{k}=\frac{k}{2}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}i\sim\mathcal{U}(0,k-1)\\ \mathrm{q}({\bm{\mu}}_{i}\mid{\bm{x}},\lambda_{i})\end{subarray}}\Big[\alpha_{i+1}\|{\bm{x}}-\hat{{\bm{x}}}_{i}\|_{2}^{2}\Big]$ (4)
@@
$\mathrm{q}({\bm{\mu}}_{i}\mid{\bm{x}},\lambda_{i})=\operatorname*{\mathbb{E}}_{\mathrm{p}({\bm{\mu}}_{0})}\big[\mathrm{p}({\bm{\mu}}_{i}\mid{\bm{\mu}}_{0},{\bm{x}},\lambda_{i})\big],\quad\hat{{\bm{x}}}_{i}=f_{{\bm{\theta}}}({\bm{\mu}}_{i},\lambda_{i})\quad\text{and}\quad\lambda_{i}=\lambda_{0}+\sum_{j=1}^{i}\alpha_{j}.$ (5)
@@
[…]
@@ 3.3 Variance Reduction
Algorithm 2 Estimating the BSI training loss
0: Initial precision $\lambda_{0}$,
precision schedule $\alpha_{i}$ for $i\in[k]$
@@ 4 Discussion
We are aware of two generative models that are closely related to BSI, BFN (Graves et al., 2023) and VDM (Kingma et al., 2023). BFNs are generative models motivated from an information theory perspective with a sender and a receiver communicating about the sample. We show in Section˜A.1 that BFNs are a special case of our framework in Section˜3 if you choose $\gamma_{0}=\infty$ instead of $\gamma_{0}=\lambda_{0}$ for the initial belief prior, meaning that sampling always starts from the deterministic ${\bm{\mu}}_{0}={\bm{0}}$. VDM are a type of DM that have shown excellent performance in likelihood-based modeling. They are similar to BSI insofar as they specify the distribution of latent variables directly rather than defining a Markovian noising process as classical DMs do.
Table 1: Central structures of VDM, BFN and BSI. To improve comparability, we parametrize VDM in terms of the SNR $\nu$. BFN and BSI are parametrized with the precision $\alpha$ as introduced in Section˜3. $\bm{\varepsilon}_{i}\sim\mathcal{N}({\bm{0}},{\bm{I}})$ is sampling noise.
@@
BFN $\mathrm{q}({\bm{\mu}}\mid{\bm{x}},\lambda)=\mathcal{N}_{\mathrm{P}}(\nicefrac{{(\lambda-\lambda_{0})}}{{\lambda}}\,{\bm{x}},\nicefrac{{\lambda^{2}}}{{(\lambda-\lambda_{0})}})$ ${\bm{\mu}}_{0}={\bm{0}}$ ${\bm{\mu}}_{i}=\frac{\lambda_{i-1}{\bm{\mu}}_{i-1}+\alpha_{i}\big(\hat{{\bm{x}}}_{i-1}+\sqrt{\frac{1}{\alpha_{i}}}\bm{\varepsilon}_{i}\big)}{\lambda_{i-1}+\alpha_{i}}$
@@
Despite this shared ELBO form, the models vary significantly. Table˜1 lists the encoding distribution $\mathrm{q}(\bm{\psi}\mid{\bm{x}},\omega)$ for each model, their prior, from which they begin the sampling process, and the update step that the models iterate during sampling. First, we see that VDM starts sampling from a standard Normal vector and BFN from the deterministic $\bm{0}$}. Only BSI allows sampling from an initial precision $\lambda_{0}$ less than $1$, which has been shown to improve sample diversity in consistency models (Song & Dhariwal, 2024). Second, the update step shared between BSI and BFN is significantly simpler than the VDM update with respect to the precision parameter and does not require working in log-space for numerical stability as recommended for VDM (Kingma et al., 2023).
For the encoding distribution $\mathrm{q}(\bm{\psi}\mid{\bm{x}},\omega)$, which provides the training inputs when the models optimize their ELBO, we turn to Fig.˜2. First, we note that BFN adds little noise overall due to their noise variance $\nicefrac{{(\lambda-\lambda_{0})}}{{\lambda^{2}}}$ going to 0 for both small and large $\lambda$. Next, we notice the encoding distribution $\mathrm{q}(\bm{\psi}\mid{\bm{x}},\underaccent{\bar}{\omega})$ with the most noise at $t=0$. While it agrees exactly with the latent prior used for sampling for BSI and BFN, for VDM it becomes approximately $\mathcal{N}_{\mathrm{P}}(0.08{\bm{x}},1)$, which differs significantly from the standard Normal prior for sampling. In fact, the image motif is still clearly discernible in the distribution mean for VDM at its maximum noise level. The amount of signal remaining in the mean for BSI at high noise levels is counteracted by much higher noise variance, e.g. $15.85$ at $t=0.1$ for BSI compared to $0.96$ for VDM.
@@ 5.1 Preconditioning
It has long been known in the context of DMs that training models to predict $\bm{x}$} directly from a noisy input can hinder learning and limit sample quality (Karras et al., 2022; Ho et al., 2020). For probabilistic modeling, it is especially important that the model prediction stays close to the true sample if the input is already at a low noise level to achieve high ELBOs. This can be seen, for example, in Corollary˜3.6 where prediction errors for high-precision input beliefs with large $\lambda$ have a higher weight. Instead of predicting $\bm{x}$}, DMs commonly either predict a variation of the noise in the model input (Ho et al., 2020; Song et al., 2021a) or an adaptive mixture of the noise and the true sample (Salimans & Ho, 2021). In the end, these approaches amount to adding a skip connection to the model with specific weights.
@@ 5.3 Hyperparameters
Apart from $f_{{\bm{\theta}}}$, BSI has three hyperparameters, $\lambda_{0}$, $\alpha_{\mathrm{M}}$ and $\alpha_{\mathrm{R}}$. $\lambda_{0}$ should be small enough that the initial belief covers the whole data distribution. We have found experimentally that $\lambda_{0}=10^{-2}$ works well for images rescaled to [1,1][-1,1]. This agrees with the finding of Song & Dhariwal (2024) that large initial noise scales improve sample diversity in consistency models.
@@ 5.4 Architecture
$\begin{pmatrix}{\bm{\mu}}&\sin(2^{i}\pi{\bm{\mu}})&\cos(2^{i}\pi{\bm{\mu}})\end{pmatrix}\in\mathbb{R}^{(3+2(n_{\mathrm{max}}-n_{\mathrm{min}}))\cdot n}\qquad i\in{n_{\mathrm{min}},\ldots,n_{\mathrm{max}}}$ (19)
@@
ViTs are a more recent architecture inspired by the success of transformers (Vaswani et al., 2017). They represent images as a set of patches with a 2D position embedding and process them with global attention, in contrast to convolutional architectures like the U-Net where communication happens primarily locally. We opt for the Diffusion Transformer (DiT) architecture (Peebles & Xie, 2023) which has been shown to improve sample quality over variants of the (Ho et al., 2020) U-Net model.
@@ 6 Experiments
We evaluate BSI on the ImageNet (Deng et al., 2009) and CIFAR10 (Krizhevsky, 2009) datasets in terms of log-likelihood and sample quality. While BSI as a method is general and not specific to images, we chose image datasets, because they are established benchmarks in the probabilistic modeling literature. In our experiments, we compare against BFN (Graves et al., 2023) and VDM (Kingma et al., 2023). BFNs are a special case of our framework (see Section˜3) and provides an important reference point for the effect of the non-deterministic hyper-prior $\mathrm{p}({\bm{\mu}}_{0})$ in BSI. VDMs are a representative of the diffusion family of models specifically designed for probabilistic modeling that is structurally similar to BSI as we explained in Section˜4.
@@
6.1 ImageNet32
Table 2: Log-likelihood in BPD and sample quality (FID) against the test set on ImageNet32. We compute standard deviations over 3 seeds. Lower is better.

### v3

@@ Abstract
We derive a novel generative model from iterative Gaussian posterior inference. By treating the generated sample as an unknown variable, we can formulate the sampling process in the language of Bayesian probability. Our model uses a sequence of prediction and posterior update steps to iteratively narrow down the unknown sample starting from a broad initial belief. In addition to a rigorous theoretical analysis, we establish a connection between our model and diffusion models and show that it includes Bayesian Flow Networks (BFNs) as a special case. In our experiments, we demonstrate that our model improves sample quality on ImageNet32 over both BFNs and the closely related Variational Diffusion Models, while achieving equivalent log-likelihoods on ImageNet32 and ImageNet64.
Machine Learning, ICML
[…]
@@ 1 Introduction
The field of deep learning has produced a multitude of generative models over the years (Harshvardhan et al., 2020). Variational autoencoders, for example, learn the data distribution by compressing data into a lower-dimensional representation (Kingma and Welling, 2013). Normalizing flows learn to map between a prior and the data distribution via invertible transformations, enabling exact likelihood computation (Rezende and Mohamed, 2015). Generative adversarial networks generate samples by pitting two models against each other such that one proposes artificial data samples while the other tries to distinguish real and generated (Goodfellow et al., 2014). Recently, diffusion models (DMs) have become a cornerstone of generative modeling (Sohl-Dickstein et al., 2015; Ho et al., 2020). They define a multi-step forward process that gradually adds noise to the data, turning it into pure noise. Then, a model is trained to reverse this process, enabling the generation of new data samples by starting from noise and iteratively denoising.
@@ 3 Sample Generation with Posterior Inference
Algorithm 1 Sampling with posterior inference
0: Initial precision $\lambda_{0}$,precision schedule $\alpha_{i}$ for $i\in[k]$
@@
Let $f_{{\bm{\theta}}}:\mathbb{R}^{n}\times\mathbb{R}_{+}\to\mathbb{R}^{n}$ be a learned model with parameters $\bm{\theta}$} that estimates which unknown sample $\bm{x}$} we have observed so far from our current belief $({\bm{\mu}}_{i},\lambda_{i})$. We estimate $\bm{x}$} as $\hat{{\bm{x}}}_{i-1}=f_{{\bm{\theta}}}({\bm{\mu}}_{i-1},\lambda_{i-1})$ and sample a noisy measurement ${\bm{y}}_{i}\sim\mathcal{N}_{\mathrm{P}}(\hat{{\bm{x}}}_{i-1},\alpha_{i})$ of $\hat{{\bm{x}}}_{i-1}$ in place of $\bm{x}$} with precision $\alpha_{i}$. Then, we can update our belief with ${\bm{y}}_{i}$ and Lemma˜2.1 to the posterior $({\bm{\mu}}_{i},\lambda_{i})$. Now, we alternate between these two steps, i.e. predicting and taking a noisy measurement followed by updating our current belief, until the posterior precision $\lambda_{i}$ is sufficient. Finally, we return $\hat{{\bm{x}}}^{*}=f_{{\bm{\theta}}}({\bm{\mu}}_{k},\lambda_{k})$ as our generated sample. See Fig.˜1 for a visual explanation and ˜9 for a formal description.
@@ 3.1 Evidence Lower Bound
By interpreting BSI as a hierarchical latent variable model, we derive an ELBO (Kingma and Welling, 2013), i.e. a lower bound on $\log\mathrm{p}({\bm{x}})$ assigned to a data point by our model. The ELBO will then serve as a natural training target for $f_{{\bm{\theta}}}$ to ensure that true data samples have high likelihood under our model.
@@
$\mathrm{p}({\bm{x}})=\operatorname*{\mathbb{E}}_{\mathrm{p}({\bm{\mu}}_{0})\cdot\mathrm{p}({\bm{\mu}}_{1}\mid{\bm{\mu}}_{0})\cdots\mathrm{p}({\bm{\mu}}_{k}\mid{\bm{\mu}}_{k-1}))}[\mathrm{p}({\bm{x}}\mid{\bm{\mu}}_{k})].$ (2)
@@
$\mathcal{L}_{\mathrm{R}}$ $=\operatorname*{\mathbb{E}}_{\mathrm{q}({\bm{\mu}}_{k}\mid{\bm{x}},\lambda_{k})}\big[{\shortminus}\negthinspace\log\mathcal{N}_{\mathrm{P}}({\bm{x}}\mid\hat{{\bm{x}}}_{k},\alpha_{\mathrm{R}})\big]$ (4)
$\mathcal{L}_{\mathrm{M}}^{k}$ $=\frac{k}{2}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}i\sim\mathcal{U}(0,k-1)\\ \mathrm{q}({\bm{\mu}}_{i}\mid{\bm{x}},\lambda_{i})\end{subarray}}\Big[\alpha_{i+1}\|{\bm{x}}-\hat{{\bm{x}}}_{i}\|_{2}^{2}\Big]$
@@
$\begin{array}[]{c}\mathrm{q}({\bm{\mu}}_{i}\mid{\bm{x}},\lambda_{i})=\operatorname*{\mathbb{E}}_{\mathrm{p}({\bm{\mu}}_{0})}\big[\mathrm{p}({\bm{\mu}}_{i}\mid{\bm{\mu}}_{0},{\bm{x}},\lambda_{i})\big],\\ \hat{{\bm{x}}}_{i}=f_{{\bm{\theta}}}({\bm{\mu}}_{i},\lambda_{i})\quad\text{and}\quad\lambda_{i}=\lambda_{0}+\sum_{j=1}^{i}\alpha_{j}.\end{array}$ (5)
@@
$\mathcal{L}_{\mathrm{R}}$ $=\operatorname*{\mathbb{E}}_{\mathrm{q}({\bm{\mu}}_{\lambda_{\mathrm{M}}}\mid{\bm{x}},\lambda_{\mathrm{M}})}\big[{\shortminus}\negthinspace\log\mathcal{N}_{\mathrm{P}}({\bm{x}}\mid\hat{{\bm{x}}}_{\lambda_{\mathrm{M}}},\alpha_{\mathrm{R}})\big]$ (6)
$\mathcal{L}_{\mathrm{M}}^{\infty}$ $=\frac{\alpha_{\mathrm{M}}}{2}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\lambda\sim\mathcal{U}(\lambda_{0},\lambda_{\mathrm{M}})\\ \mathrm{q}({\bm{\mu}}_{\lambda}\mid{\bm{x}},\lambda)\end{subarray}}\big[\|{\bm{x}}-\hat{{\bm{x}}}_{\lambda}\|_{2}^{2}\big]$
@@ 3.3 Variance Reduction
[…]
3.4 Training & Sampling
Algorithm 2 Estimating the BSI training loss
@@ 3.4 Training & Sampling
Algorithm 3 Sampling with BSI
0: Initial precision $\lambda_{0}$,precision schedule $\alpha_{i}$ for $i\in[k]$
@@ 4 Discussion
We are aware of two generative models that are closely related to BSI, BFN (Graves et al., 2023) and VDM (Kingma et al., 2023). BFNs are generative models motivated from an information theory perspective with a sender and a receiver communicating about the sample. As we show in Section˜A.1, BFNs are a special case of our framework in Section˜3 if we translate them to the probabilistic perspective. They correspond to choosing $\gamma_{0}=\infty$ and $\lambda_{0}=1$, meaning that sampling always starts from the deterministic belief $({\bm{\mu}}_{0},\lambda_{0})=({\bm{0}},1)$. In contrast, BSI chooses $\gamma_{0}=\lambda_{0}$, i.e. the noise in the initial belief corresponds to our confidence in it, and leaves $\lambda_{0}$ as a hyperparameter, which we investigate in Section˜6. VDM are a type of DM that have shown excellent performance in likelihood-based modeling. They are similar to BSI insofar as they specify the distribution of latent variables directly rather than defining a Markovian noising process as classical DMs do.
Table 1: Central structures of VDM, BFN and BSI. To improve comparability, we parametrize VDM in terms of the SNR $\nu$. BFN and BSI are parametrized with the belief precision $\lambda$ as introduced in Section˜3. $\bm{\varepsilon}_{i}\sim\mathcal{N}({\bm{0}},{\bm{I}})$ is sampling noise.
@@
BFN $\mathrm{q}({\bm{\mu}}\mid{\bm{x}},\lambda)=\mathcal{N}_{\mathrm{P}}(\nicefrac{{(\lambda-1)}}{{\lambda}}\,{\bm{x}},\nicefrac{{\lambda^{2}}}{{(\lambda-1)}})$ ${\bm{\mu}}_{0}={\bm{0}}$ ${\bm{\mu}}_{i}=\frac{\lambda_{i-1}{\bm{\mu}}_{i-1}+\alpha_{i}\big(\hat{{\bm{x}}}_{i-1}+\sqrt{\frac{1}{\alpha_{i}}}\bm{\varepsilon}_{i}\big)}{\lambda_{i-1}+\alpha_{i}}$
@@
Despite this shared ELBO form, the models vary significantly. Table˜1 lists the encoding distribution $\mathrm{q}(\bm{\psi}\mid{\bm{x}},\omega)$ for each model, their prior, from which they begin the sampling process, and the update step that the models iterate during sampling. First, we see that VDM starts sampling from a standard Normal vector and BFN from the deterministic $\bm{0}$}. Only BSI allows sampling from an initial precision $\lambda_{0}$ less than $1$, which has been shown to improve sample diversity in consistency models (Song and Dhariwal, 2024). Second, the update step shared between BSI and BFN is significantly simpler than the VDM update with respect to the precision parameter and does not require evaluation in log-space for numerical stability as recommended for VDM (Kingma et al., 2023).
For the encoding distribution $\mathrm{q}(\bm{\psi}\mid{\bm{x}},\omega)$, which provides the training inputs when the models optimize their ELBO, we turn to Fig.˜2. First, we note that BFN adds little noise overall due to their noise variance $\nicefrac{{(\lambda-1)}}{{\lambda^{2}}}$ going to 0 for both small and large $\lambda$. Next, we notice the encoding distribution $\mathrm{q}(\bm{\psi}\mid{\bm{x}},\underaccent{\bar}{\omega})$ with the most noise at $t=0$. While it agrees exactly with the latent prior used for sampling for BSI and BFN, for VDM it becomes approximately $\mathcal{N}_{\mathrm{P}}(0.08{\bm{x}},1)$, which differs significantly from the standard Normal prior for sampling. In fact, the image motif is still clearly discernible in the distribution mean for VDM at its maximum noise level. The amount of signal remaining in the mean for BSI at high noise levels is counteracted by much higher noise variance, e.g. $15.85$ at $t=0.1$ for BSI compared to $0.96$ for VDM.
@@ 5.1 Preconditioning
It has long been known in the context of DMs that training models to predict $\bm{x}$} directly from a noisy input can hinder learning and limit sample quality (Karras et al., 2022; Ho et al., 2020). For probabilistic modeling, it is especially important that the model prediction stays close to the true sample if the input is already at a low noise level to achieve high ELBOs. This can be seen, for example, in Corollary˜3.6 where prediction errors for high-precision input beliefs with large $\lambda$ have a higher weight. Instead of predicting $\bm{x}$}, DMs commonly either predict a variation of the noise in the model input (Ho et al., 2020; Song et al., 2021a) or an adaptive mixture of the noise and the true sample (Salimans and Ho, 2021). In the end, these approaches amount to adding a skip connection to the model with specific weights.
@@ 5.3 Hyperparameters
Apart from $f_{{\bm{\theta}}}$, BSI has three hyperparameters, $\lambda_{0}$, $\alpha_{\mathrm{M}}$ and $\alpha_{\mathrm{R}}$. $\lambda_{0}$ should be small enough that the initial belief covers the whole data distribution. We have found experimentally that $\lambda_{0}=10^{-2}$ optimizes likelihoods and sample quality at the same time for images rescaled to [1,1][-1,1], see Section˜6.3. This agrees with the finding of Song and Dhariwal (2024) that large initial noise scales improve sample diversity in consistency models.
@@ 5.4 Architecture
$\begin{pmatrix}{\bm{\mu}}&\sin(2^{i}\pi{\bm{\mu}})&\cos(2^{i}\pi{\bm{\mu}})\end{pmatrix}\ i\in{n_{\mathrm{min}},\ldots,n_{\mathrm{max}}}$ (19)
@@
ViTs are a more recent architecture inspired by the success of transformers (Vaswani et al., 2017). They represent images as a set of patches with a 2D position embedding and process them with global attention, in contrast to convolutional architectures like the U-Net where communication happens primarily locally. We opt for the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023) which has been shown to improve sample quality over variants of the (Ho et al., 2020) U-Net model.
@@ 6 Experiments
We evaluate BSI on the ImageNet (Deng et al., 2009) dataset in terms of log-likelihood and sample quality and on CIFAR10 (Krizhevsky, 2009) in terms of log-likelihood. While BSI as a method is general and not specific to images, we chose image datasets, because they are established benchmarks in the probabilistic modeling literature. In our experiments, we compare against BFN (Graves et al., 2023) and VDM (Kingma et al., 2023). BFNs are a special case of our framework (see Section˜3) and provides an important reference point for the effect of the non-deterministic hyper-prior $\mathrm{p}({\bm{\mu}}_{0})$ in BSI. VDMs are a representative of the diffusion family of models specifically designed for probabilistic modeling that is structurally similar to BSI as we explained in Section˜4.
@@
6.1 ImageNet
Table 2: Log-likelihood in BPD and sample quality (FID) against the test set on ImageNet. We compute standard deviations over 3 seeds.